OpenStudy (anonymous):

@FibonacciChick666 Minimum value of \[sqrt{10x^2+12x+6}\]

OpenStudy (fibonaccichick666):

\[ \)

OpenStudy (fibonaccichick666):

\[\sqrt{10x^2+12x+6}\]

OpenStudy (anonymous):

For what value of x does f(x) take on a minimum if [...]

OpenStudy (anonymous):

What is the minimum value of f(x) if [...]

OpenStudy (fibonaccichick666):

Let x=\(\frac{3}{5}\)

OpenStudy (fibonaccichick666):

\[\sqrt{10(\frac{-3}{5})^2+12(\frac{-3}{5})+6}\]